# Blume-Capel model analysis with microcanonical population annealing method

### v2

@@ III.2 Floor microcanonical procedure
Figure 2: Example entropy estimate for Blume-Capel model with $D=1.5$ and linear system size $L=32$. The orange points are calculated using the ceiling algorithm, the blue points are calculated using the floor algorithm, the right red vertical line marks the rightmost value of $S^{c}(E)$, the left red vertical line marks the leftmost value of $S^{f}(E)$. The inner green vertical lines mark the stitching region. Details are given in the text.
@@ III.4 Stitching the parts of entropy together
$S^{c}(E)$ $=$ $\ln(\epsilon(E))+\sum_{E>E_{c}}\ln(1-\epsilon(E^{\prime})),$ (4)
$S^{f}(E)$ $=$ $\ln(\epsilon(E))+\sum_{E<E_{f}}\ln(1-\epsilon(E^{\prime})).$ (5)
@@ IV Algorithm realization
Figure 5: Variation of the relative DoS error $\delta R$ on the number of replicas $R$. The dashed line shows the slop proportional to $1/R^{1/2}$. 2D square lattice Ising model with $L{=}20$.
@@ V Simulation results
Figure 6: The entropy $S(E)$ is calculated for the ratios $\Delta{/}J{=}0$ and 1.5 in the top row and $\Delta{/}J{=}1.966$ and 1.99 in the bottom row. The number of replicas is $72^{1}$ 7 and the system size is $L=64$.
@@ V.1 Specific heat analysis
Figure 8: Maximum value of the specific heat of Blume-Cappel model. Red circles - $\Delta/J{=}1.95$, black squares - $\Delta/J{=}1.96$,purple stars - $\Delta/J{=}1.962$, magenta circles - $\Delta/J{=}1.966$.

### v3

# Blume-Capel model analysis with a microcanonical population annealing method

@@ III.2 Floor microcanonical procedure
Figure 2: Example entropy estimate for Blume-Capel model with $D=1.5$ and linear system size $L=32$. The blue solid squares are calculated using the ceiling algorithm, the red open circles are calculated using the floor algorithm, the right dashed red vertical line marks the rightmost value of $S^{c}(E)$, the left dashed red vertical line marks the leftmost value of $S^{f}(E)$. The inner green solid vertical lines mark the stitching region. Details are given in the text.
@@ III.4 Stitching the parts of entropy together
$S^{c}(E)$ $=$ $\ln(\epsilon(E))+\sum_{E^{\prime}>E}\ln(1-\epsilon(E^{\prime})),$ (4)
$S^{f}(E)$ $=$ $\ln(\epsilon(E))+\sum_{E^{\prime}<E}\ln(1-\epsilon(E^{\prime})).$ (5)
@@ IV Algorithm realization
Figure 5: Variation of the relative DoS error $\delta g$ on the number of replicas $R$. The dashed line shows the slop proportional to $1/R^{1/2}$. 2D square lattice Ising model with $L{=}20$.
@@ V Simulation results
Figure 6: The entropy $S(E)$ is calculated for the ratios $\Delta{/}J{=}0$ and 1.5 in the top row and $\Delta{/}J{=}1.966$ and 1.99 in the bottom row. The number of replicas is $2^{17}$ and the system size is $L=64$.
@@ V.1 Specific heat analysis
Figure 8: Maximum value of the specific heat of Blume-Cappel model. Red circles - $\Delta/J{=}1.95$, black squares - $\Delta/J{=}1.96$,purple stars - $\Delta/J{=}1.962$, magenta pentagons - $\Delta/J{=}1.966$.